OpenStudy (ckenn):

Write an equation of a line that passes through the point (3, 2) and is parallel to the line y = 3x − 4. a) y = 3x + 7 b) y = 3x - 7 c) y = 1/3x + 2 d) y = 1/3x - 2

OpenStudy (ckenn):

@zepdrix @surjithayer do you think you could help me?

zepdrix (zepdrix):

parallel means they have exactly the same slope. are you able to identify the slope of this equation? y=3x-4

OpenStudy (ckenn):

3?

zepdrix (zepdrix):

Yes, and the slope-intercept form of a straight line is given by: y=mx+b So, filling in our slope value, we have this so far: \(\large\rm y=3x+b\) That at least helps to narrow down our choice.

zepdrix (zepdrix):

Now you need to plug in the point (3,2), and determine the missing "b" value.

OpenStudy (ckenn):

thank you! so it would be "A"

OpenStudy (ckenn):

@zepdrix ?

zepdrix (zepdrix):

\[\large\rm 3=3(2)+b\]You got b=7 ? Hmm that doesn't seem quite right.

zepdrix (zepdrix):

Woops I plugged those in backwards lolol :)

zepdrix (zepdrix):

\[\large\rm 2=3(3)+b\]But again, b=7 doesn't seem quite right :)

OpenStudy (ckenn):

b = 2, is that right?

zepdrix (zepdrix):

\[\large\rm 2=3(3)+b\]\[\large\rm 2=9+b\]Subtracting 9 from each side gives\[\large\rm 2-9=b\]